# Mind-to-Image: Projecting Visual Mental Imagination of the Brain from fMRI

### v3

@@ Abstract
The reconstruction of images observed by subjects from fMRI data collected during visual stimuli has made significant strides in the past decade, thanks to the availability of extensive fMRI datasets and advancements in generative models for image generation. However, the application of visual reconstruction has remained limited. Reconstructing visual imagination presents a greater challenge, with potentially revolutionary applications ranging from aiding individuals with disabilities to verifying witness accounts in court. The primary hurdles in this field are the absence of data collection protocols for visual imagery and the lack of datasets on the subject. Traditionally, fMRI-to-image relies on data collected from subjects exposed to visual stimuli, which poses issues for generating visual imagery based on the difference of brain activity between visual stimulation and visual imagery. For the first time, we have compiled a substantial dataset (around 6h of scans) on visual imagery along with a proposed data collection protocol. We then train a modified version of an fMRI-to-image model and demonstrate the feasibility of reconstructing images from two modes of imagination: from memory and from pure imagination. This marks an important step towards creating a technology that allow direct reconstruction of visual imagery.
@@ 1 Introduction
Figure 1: Overview of our method. Left: classic fMRI-to-Image approach where a subject sees images in an MRI scanner, bold data are recorded then used to train a model that convert brain data to match the seen images. Right: in our approach we let the subject imagine previously seen images from a dataset of surrealist images (face portraits and nature landscapes) which we create. The brain data associated to the recollection of such images is gathered. It is then used to train a model to reconstruct previously seen images from visual imagination brain data. At inference time we collect brain data where the subject completely imagines new images based on instructions and use the trained model of the first step to generate the imagined images.
@@
In the rapidly evolving field of neuroscience and artificial intelligence, the quest to decode the complex mechanisms of the human brain has led to groundbreaking developments, particularly in the area of functional magnetic resonance imaging (fMRI). Among the myriad applications of fMRI, one of the most intriguing has been the reconstruction of visual experiences. Traditional methods have successfully reconstructed images based on fMRI BOLD (blood-oxygen-level-dependent) data captured while subjects view specific images (Chen et al., 2022, 2023; Scotti et al., 2023). This process involves preprocessing the BOLD data to isolate voxels associated with the brain’s visual areas, generating clip embeddings from these voxels, and then employing image generation algorithms that utilize these embeddings to recreate the observed images. Such methodologies have shown promising results, especially when employing datasets like the NSD (Natural Scenes Dataset) (Allen et al., 2022), a substantial 7T fMRI dataset designed to bridge cognitive neuroscience and artificial intelligence. This dataset, featuring extensive scans from eight subjects, has become a cornerstone for research in this domain.
@@
The intersection of cognitive neuroscience and artificial intelligence holds the potential not only to advance our understanding of brain mechanisms but also to develop practical applications, such as real-time clinical tools and brain-computer interfaces. These applications are predicated on the ability to accurately interpret and reconstruct cognitive states and environmental inputs from brain activity. However, despite the progress in visual reconstruction from observed images, a significant challenge remains in the domain of visual imagination. Visual imagination encompasses both the recall of previously seen images (which we refer to as weak imagination in the paper) and the generation of entirely new images within the mind’s eye (which we refer to as strong imagination). The ability to decode and reconstruct these imagined visuals offers promising possibilities for aiding individuals with disabilities, enhancing creative processes, and even providing novel forms of evidence in legal contexts.
@@
Leveraging the MindEye model (Scotti et al., 2023), a state-of-the-art fMRI-to-image framework, we adapt its architecture to accommodate the broader dimensional complexity inherent in visual imagination data. Our approach is summarised in Fig.1. Our results, while preliminary, indicate the model’s capability to distinguish between and accurately generate portraits and landscapes based on the nature of the imagination. Although the fidelity of the reconstructions to the imagined content varies, our findings underscore the potential of this approach. With further refinement and expansion of the dataset, this research could pave the way for more precise reconstructions of visual imagination, opening new avenues for understanding and exploiting the brain’s creative and cognitive processes.
@@ 2 Related work
More recently, (Naselaris et al., 2015) developed a voxel-wise encoding model capable of decoding mental images of remembered scenes, demonstrating the relationship between memory, imagination, and visual perception. (VanRullen & Reddy, 2019) employs a combination of variational auto-encoders (VAE) and Generative Adversarial Networks (GAN) to perform robust pairwise decoding of faces with a high degree of accuracy but also to classify gender and decode which face was imagined by subjects. (Goebel et al., 2022) showed the feasibility of reading imagined letter shapes from the mind’s eye using real-time 7 Tesla fMRI, presenting a novel approach to understanding how specific visual information can be decoded from brain activity.
These studies collectively underscore the potential of fMRI data in decoding not just seen images but also those that are imagined or remembered.
3 fMRI data collection protocol for visual imagery
The development of a comprehensive data collection protocol is one the main contributions of our study, aimed at capturing the complex neural correlates of visual imagination. Drawing inspiration from the NSD paper (Allen et al., 2022), our protocol is designed to explore two distinct facets of visual imagination: weak imagination (imagination from memory) and strong imagination (pure imagination). This dual approach enables us to study how the brain navigates between recalling visual information and generating novel visual constructs.
@@ 3.2 Weak imagination: data collection protocol
The protocol for weak imagination aims at recording brain activity corresponding to the recall of previously seen images. In this phase, subjects are first shown an image for 3 seconds, followed by a 1-second rest period. This sequence is repeated with three different images, each selected from our curated set of 600 surrealist portraits and 600 surrealist landscapes. To engage the brain’s imaginative capacities, the same images are then flashed for 0.1 seconds, prompting the subject to imagine the image for 5 seconds before resting for another second. This process is repeated over a span of 6 hours, encompassing a total of 1200 images. Out of the data collected, 100 data points (comprising brain activity and the corresponding image) are earmarked for validation and evaluation, with the remainder serving as the training set for our fMRI-to-image model. The evaluation focuses on the model’s ability to generate images that not only match the correct modality (face portrait or nature landscape) but also reflect the contents of the original image.
@@ 3.3 Strong imagination: data collection protocol
For strong imagination, the process pivots towards the generation of entirely new visual content based on verbal prompts. Subjects are instructed to imagine a portrait or landscape evoking one of ten primal emotions (such as fear, happiness, love, etc.). This task is designed to stimulate the pure imagination process, encouraging the brain to construct novel visual imagery without the reference of previously seen images. Each imagination session lasts for 6 seconds, followed by a 1-second rest, cycling through combinations of landscape/portrait and emotional prompts. This cycle is repeated 10 times. Unlike the data collected for weak imagination, this dataset is solely used during the model’s inference stage. The evaluation here is similar to the weak imagination evaluation: quantitively assess if the model matches the correct modality (face portrait or nature landscape) and qualitatively assess if the contents of the generated image matches the visual imagery of the subject. To ensure that we recall what the subject has imagined, the subject provides an oral description of the visual contents of the imagined image.
@@ 3.4 Pre-processing of raw BOLD data and masking of the regions of interest
After the raw BOLD (blood-oxygen-level-dependent) data is collected, we apply a preprocessing routine to provide the model with data that decorrelates all the noise that impair the following steps. We first co-register all BOLD data with a T1 anatomical scan. Then, using a General Linear Model (GLM) with the Glover hemodynamic response function (HRF) (implemented via the Nilearn library (Abraham et al., 2014)), we extracted beta values associated with each voxel for every distinct event, including both the displayed images and the periods of imagination. This allwos us to isolate the specific brain activity corresponding to each event, which is common practice for such tasks.
In the subsequent data pre-processing phase, we identify the brain regions most instrumental to visual imagination and encoding, using a GLM model to manually discern these areas of interest. Unlike the NSD dataset, which predominantly focuses on visual areas, our study necessitated the creation of a custom mask to encapsulate the regions identified as crucial in our specific context of weak and strong imagination. This aforementioned mask was manually curated based on the GLM analysis results, ensuring that only voxels corresponding to our defined regions of interest were included in further analyses. The delineation and analysis of these regions, and how they diverge from the predominantly visual areas highlighted in the NSD study, is studied in the results section.
@@
4 Training the fMRI-to-Image model
We start from the state-of-the-art MindEye model which reconstructs images viewed by subjects from fMRI data by combining two pipelines: a high-level (semantic) pipeline and a low-level (perceptual) pipeline, both grounded in a residual MLP backbone structure. The high-level pipeline maps fMRI voxels to the CLIP Vision Transformer L/14 image space, capturing the semantics of the image. In contrast, the low-level pipeline targets the perceptual fidelity of reconstructed images by mapping voxels to a Variational Autoencoder’s (VAE) image embedding space. They leverage contrastive and mean squared error (MSE) losses for training of its projection and reconstruction capabilities. Techniques such as MixCo are employed to improve classification and retrieval performance by generating synthetic training data.
4.1 Adapting MindEye to larger masks for visual imagery
To accommodate the increased dimensionality inherent in our fMRI BOLD voxel data, which covers extensive brain areas notably larger than only the visual areas of the brain, we adapted MindEye’s architecture. This adjustment aimed to manage the significant growth in input dimensions without exponentially increasing the model’s parameter count. We thus simply refine the architecture of the MLP so that the model efficiently processes the data while maintaining high performance without overburdening computational resources.
4.2 Experiments
Our experimental procedures are as outlined below:
@@ 4.2 Experiments
  • We finally freeze the parameters of the fMRI-to-Image model. This frozen model is then employed to perform inference on the dataset derived from the strong imagination data collection protocol, with the goal of visualising the purely imagined images.
@@ 5.1 What happens in the brain during visualisation, weak and strong imagination?
The results of the first test (A) show that the brain regions mainly involved in normal visual perception are, as expected, the primary and secondary visual areas in the occipital lobe, the superior parietal lobule and the bottom of the temporal lobe.
@@ 5.2 Can we reconstruct images from weak imagination / memory?
In the exploration of weak imagination, where subjects were prompted to recall previously seen images, our results highlight the capacity of the model to identify to the original class of the images. Specifically, if a subject recalled a face portrait, the reconstructed image is almost always a face portrait, and similarly for nature landscapes. This observation suggests a the ability of the model to grasp the category of the imagined content accurately within the BOLD data. To quantitatively assess this capability, we fine-tune a pretrained ResNet50 model (He et al., 2016) to classify images as either face portraits or nature landscapes using the images in our dataset. We apply this classifier to the reconstructed images from the weak imagination tasks of the validation set. We obtain a classification accuracy of 91% in correctly identifying the reconstructed images as belonging to their original categories of either portraits or landscapes. This level of accuracy demonstrates the model’s ability in capturing the category information within the neural activity during weak imagination.
@@
While the model demonstrated a robust ability to regenerate the broad category of the original images, the fidelity of content reconstruction within these images presented a more varied outcome. Some reconstructions were close to the remembered images. However, others diverged more significantly, indicating a variability in the model’s ability to precisely capture and reconstruct the nuanced details of each imagined scene. This variance in reconstruction fidelity can be attributed, in part, to the limitations imposed by the dataset’s size and the duration of the fMRI scanning sessions. With only 6 hours of scans and 1200 images constituting the dataset, the goal for training a model from scratch to achieve perfect reconstruction is out of reach. Nevertheless, these initial results are promising, suggesting a capacity for accurate content reproduction. It is reasonable to anticipate that with an expanded dataset, the fidelity of the reconstructed images to their original counterparts would see significant improvement, enhancing both the precision and consistency of the content reproduction in future iterations.

### v4

@@ Abstract
The reconstruction of images observed by subjects from fMRI data collected during visual stimuli has made strong progress in the past decade, thanks to the availability of extensive fMRI datasets and advancements in generative models for image generation. However, the application of visual reconstruction has remained limited. Reconstructing visual imagination presents a greater challenge, with potentially revolutionary applications ranging from aiding individuals with disabilities to verifying witness accounts in court. The primary hurdles in this field are the absence of data collection protocols for visual imagery and the lack of datasets on the subject. Traditionally, fMRI-to-image relies on data collected from subjects exposed to visual stimuli, which poses issues for generating visual imagery based on the difference of brain activity between visual stimulation and visual imagery. For the first time, we have compiled a substantial dataset (around 6h of scans) on visual imagery along with a proposed data collection protocol. We then train a modified version of an fMRI-to-image model and demonstrate the feasibility of reconstructing images from two modes of imagination: from memory and from pure imagination. The resulting pipeline we call Mind-to-Image marks a step towards creating a technology that allow direct reconstruction of visual imagery.
@@ 1 Introduction
Figure 1: Overview of our Mind-to-Image pipeline which allows to reconstruct images from mental imagery. Left: overview of MindEye, the fMRI-to-Image model used in our Mind-to-Image approach. The stimuli (an image seen or in our case imagined) and the associated fMRI data is fed to a high-level pipeline (composed of an MLP and Diffusion Prior) and a low-level pipeline to create aligned CLIP embeddings of the fMRI data. Middle: fMRI data collection protocols. In the classic fMRI-to-Image approach, a subject sees images in an MRI scanner, BOLD (Blood-Oxygen Level Dependent) data are recorded then used to train an fMRI-to-Image model that convert brain data to match the seen images. In our approach, we devise two mental imagery protocols: weak and strong imagination. For weak imagination, we let the subject imagine previously seen images from a dataset of surrealist images (face portraits and nature landscapes) which we create. The brain data associated to the recollection of such images is gathered. For strong imagination, we collect brain data where the subject completely imagines new images based on instructions. Right: our Mind-to-Image pipeline. We use the weak imagination fMRI data to train an fMRI-to-Image model to reconstruct recollected images from visual imagination brain data. At inference time we use strong imagination fMRI data and this trained model to generate the imagined images, thanks to transfer learning.
@@
In the rapidly evolving field of neuroscience and artificial intelligence, the quest to decode the complex mechanisms of the human brain has led to groundbreaking developments, particularly in the area of functional magnetic resonance imaging (fMRI). Among the myriad applications of fMRI, one of the most intriguing has been the reconstruction of visual experiences. Traditional methods have successfully reconstructed images based on fMRI BOLD (blood-oxygen-level-dependent) data captured while subjects view specific images (Chen et al., 2022, 2023; Scotti et al., 2023). This process involves preprocessing the BOLD data to isolate voxels associated with the brain’s visual areas, generating clip embeddings from these voxels, and then employing image generation algorithms that utilize these embeddings to recreate the observed images. Such methodologies have shown promising results, especially when employing datasets like the NSD (Natural Scenes Dataset) (Allen et al., 2022), a substantial 7T fMRI dataset designed to bridge cognitive neuroscience and artificial intelligence. This dataset, featuring extensive scans from eight subjects, has had tremendous impact for research in this domain.
@@
Progress in cognitive neuroscience and artificial intelligence has not only allowed to advance our understanding of brain mechanisms but also to holds the promise to develop practical applications, such as real-time clinical tools and brain-computer interfaces. However, despite the progress in visual reconstruction from observed images, a challenge remains in the domain of visual imagination. Visual imagination encompasses both the recall of previously seen images (which we refer to as weak imagination in the paper) and the generation of entirely new images within the mind’s eye (which we refer to as strong imagination). The ability to decode and reconstruct these imagined visuals offers a wide range of promising possibilities from aiding individuals with disabilities, to enhancing creative processes.
@@
Leveraging the MindEye model (Scotti et al., 2023), a state-of-the-art fMRI-to-image framework, we adapt its architecture to adapt to the larger dimensional complexity inherent in visual imagination data. Our approach is summarised in Fig.1. Our results, while preliminary, indicate the model’s capability to, at least to some degree, generate portraits and landscapes based on the nature of the imagination. Although the fidelity of the reconstructions to the imagined content varies, which is measured quantitatively through usual metrics in the field, our findings underscore the potential of this approach. With further refinement and expansion of the dataset, this research could pave the way for more precise reconstructions of visual imagination, with the promise of a better understanding of the brain’s creative and cognitive processes.
@@ 2 Related work
More recently, (Naselaris et al., 2015) developed a voxel-wise encoding model capable of decoding mental images of remembered scenes, demonstrating the relationship between memory, imagination, and visual perception. (VanRullen & Reddy, 2019) employs a combination of variational auto-encoders (VAE) and Generative Adversarial Networks (GAN) to perform pairwise decoding of faces with a high degree of accuracy but also to classify gender and decode which face was imagined by subjects. (Goebel et al., 2022) showed the feasibility of reading imagined letter shapes from the mind’s eye using real-time 7 Tesla fMRI, presenting a novel approach to understanding how specific visual information can be decoded from brain activity.
These studies collectively show the potential of fMRI data in decoding not just seen images but also those that are imagined or remembered.
More close to our work, a recent paper (Koide-Majima et al., 2024) provides preliminary results on reconstructing images from memory. While their results are encouraging, their reconstruction pipeline backbone is not state-of-the-art like MindEye is, which partly explains the results obtained.
3 fMRI data collection for visual imagery
The development of a comprehensive data collection protocol is one the main contributions of our study, aimed at capturing the complex neural correlates of visual imagination. Drawing inspiration from the NSD paper (Allen et al., 2022), our protocol is designed to explore two distinct modes of visual imagination: weak imagination (imagination from memory) and strong imagination (pure imagination). This approach enables us to study how the brain navigates between recalling visual information and generating novel visual constructs.
@@ 3.1 Image dataset used: Surrealism
3.2 fMRI parameters
Data were collected on Siemens 3T PrismaFit using a 64 channels headcoil. BOLD data were acquired using the CMRR multiband EPI sequence. The main parameters are: 2mm isotropic voxels with full brain coverage, TR/TE = 1300ms/27ms, MultiBand=4, partialFourier=7/8, FlipAngle=68°, P>>much-greater-than>>> >A phase encoding direction, EchoSpacing at minimum.
@@ 3.3 Weak imagination: data collection protocol
The protocol for weak imagination aims at recording brain activity corresponding to the recall of previously seen images. In this phase, subjects are first shown an image for 3 seconds, followed by a 1-second rest period. This sequence is repeated with three different images, each selected from our curated set of 600 surrealist portraits and 600 surrealist landscapes. To engage the brain’s imagination, the same images are then flashed for 0.1 seconds, prompting the subject to imagine the image for 5 seconds before resting for another second. This process is repeated over a span of 6 hours, encompassing a total of 1200 images. Out of the data collected, 75 data points (comprising brain activity and the corresponding image) are reserved for validation and evaluation, with the remainder serving as the training set for our fMRI-to-image model. The evaluation focuses on the model’s ability to generate images that not only match the correct modality (face portrait or nature landscape) but also reflect the contents of the original image.
@@ 3.4 Strong imagination: data collection protocol
For strong imagination, the subject is now asked to generate entirely new visual content based on verbal prompts. The subject is instructed to imagine a portrait or landscape evoking one of ten primal emotions (such as fear, happiness, love, etc.). This task is designed to stimulate the pure imagination process, encouraging the brain to construct novel visual imagery without the reference of previously seen images. Each imagination session lasts for 6 seconds, followed by a 1-second rest, cycling through combinations of landscape/portrait and emotional prompts. This cycle is repeated 10 times. Unlike the data collected for weak imagination, this dataset is solely used during the model’s inference stage. The evaluation here is similar to the weak imagination evaluation: quantitively assess if the model matches the correct modality (face portrait or nature landscape) and qualitatively assess if the contents of the generated image matches the visual imagery of the subject. However we cannot compute all the quantitative metrics as the ground truth image does not actually exists outside the subject’s mind. To ensure that we recall what the subject has imagined, the subject provides an oral description of the visual contents of the imagined image.
@@ 3.5 Pre-processing of raw BOLD data and masking of the regions of interest
After the raw BOLD (blood-oxygen-level-dependent) data is collected, we apply a preprocessing routine to provide the model with data that decorrelates all the noise that impair the following steps. We first co-register all BOLD data with a T1 anatomical scan. Then, using a General Linear Model (GLM) with the Glover hemodynamic response function (HRF) (implemented via the Nilearn library (Abraham et al., 2014)), we extracted beta values associated with each voxel for every distinct event, including both the displayed images and the periods of imagination. This allows us to isolate the specific brain activity corresponding to each event, which is common practice for such tasks.
In the subsequent data pre-processing phase, we identify the brain regions most important to visual imagination and encoding, using a GLM model to manually extract these areas of interest. Unlike the NSD dataset, which predominantly focuses on visual areas, our study required the creation of a custom mask to englobe the regions identified as crucial in our context of weak and strong imagination. This aforementioned mask was manually curated based on the GLM analysis results, ensuring that only voxels corresponding to our defined regions of interest were included in further analyses. The delineation and analysis of these regions, and how they diverge from the predominantly visual areas highlighted in the NSD study, is studied in the results section.
@@
4 Mind-to-Image pipeline
We know describe the fMRI-to-Image model used for our experiments and then outline our Mind-to-Image pipeline by connecting the dots between the data collection protocol and the fMRI-to-Image model.
4.1 fMRI-to-Image model
We start from the state-of-the-art MindEye model which reconstructs images viewed by subjects from fMRI data by combining two pipelines: a high-level (semantic) pipeline and a low-level (perceptual) pipeline. The high-level pipeline maps fMRI voxels to the CLIP Vision Transformer L/14 image space, capturing the semantics of the image. The low-level pipeline targets the perceptual fidelity of reconstructed images by mapping voxels to Stable Diffusion’s Variational Autoencoder (VAE) embedding space. They leverage contrastive and mean squared error (MSE) losses for training of its projection and reconstruction capabilities. At inference time, fMRI data associated to an image is fed through the model which outputs a low-level reconstructed image and aligned CLIP embeddings. These embeddings are fed to an UnCLIP models, which is a model able to reconstruct an image based on its CLIP embeddings. In this case, the embeddings as well as the low-level reconstructed image is fed to a Versatile Diffusion Image Variations pipeline to obtain the final reconstruction, as it is done in the original MindEye paper. The whole fMRI-to-Image pipeline is illustrated in Fig.1.
Adapting fMRI-to-Image to larger masks for visual imagery. To accommodate the increased dimensionality inherent in our fMRI BOLD voxel data, which covers extensive brain areas notably larger than only the visual areas of the brain, we adapted MindEye’s architecture. This adjustment aimed to manage the growth in input dimensions without exponentially increasing the model’s parameter count. We thus simply refine the architecture of the MLP so that the model efficiently processes the data while maintaining a reasonable cost in GPU memory.
4.2 Mind-to-Image pipeline
Our Mind-to-Image experimental procedures are as outlined below:
@@ 4.2 Mind-to-Image pipeline
  • We finally freeze the parameters of the fMRI-to-Image model. This frozen model is then employed to perform inference on the dataset derived from the strong imagination data collection protocol, with the goal of visualising the purely imagined images thanks to transfer learning.
@@ 5.1 What happens in the brain during visualisation, weak and strong imagination?
The results of the first test (A) show that the brain regions involved in normal visual perception are, as expected, the primary and secondary visual areas in the occipital lobe, the superior parietal lobule and the bottom of the temporal lobe.
@@ 5.2 Can we reconstruct images from weak imagination / memory?
Method Low-Level High-Level
PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
[…]
Table 1: Quantitative comparison of Mind-to-Image reconstruction performance against other models trained on different datasets, computed on usual metrics in the field. The results are not directly comparable, but are given to provide an element of comparison on the ranges of the values of the metrics. PixCorr is the pixel-level correlation of reconstructed and groundtruth images. SSIM (Wang et al., 2004) is the structural similarity index metric. AlexNet(2) and AlexNet(5) are the 2-way comparisons of the second and fifth layers of AlexNet (Krizhevsky et al., 2012), respectively. Inception is the 2-way comparison of the last pooling layer of InceptionV3 (Szegedy et al., 2016). CLIP is the 2-way comparison of the output layer of the CLIP-Vision model (Radford et al., 2021). EffNet-B and SwAV are distance metrics gathered from EfficientNet-B (Tan & Le, 2019) and SwAV-ResNet (Caron et al., 2020) models, respectively.
@@
In the exploration of weak imagination, where the subject is asked to recall previously seen images, our results show the capacity of the model to identify to the original class of the images. Specifically, if a subject recalled a face portrait, the reconstructed image is almost always a face portrait, and similarly for nature landscapes. This observation suggests a the ability of the model to grasp the category of the imagined content accurately within the BOLD data.
First, to quantitatively assess this capability, we fine-tune a pretrained ResNet50 model (He et al., 2016) to classify images as either face portraits or nature landscapes using the images in our dataset. We apply this classifier to the reconstructed images from the weak imagination tasks of the validation set. We obtain a classification accuracy of 91% in correctly identifying the reconstructed images as belonging to their original categories of either portraits or landscapes. This level of accuracy demonstrates the model’s ability in capturing the category information within the neural activity during weak imagination.
Then, we provide in Tab.1 the usual metrics to quantitatively assess the performance of reconstruction from weak imagination. We compare our results with MindEye (Scotti et al., 2023) Note that the comparisons are not apples-to-apples: (Scotti et al., 2023) provide results for visual stimuli experiments (not for weak imagination) on a large-scale (approx. 10x size) dataset that uses another image domain (realistic images), explaining the big gap in performance. Our results, while inferior to those of reconstructions from visual stimuli, show that we are able to have significantly better than chance results on reconstruction from weak imagination. We noted two main outliers on our quantitative results: we get poor results on the SSIM metric, and excellent results on the SwAV metric. We posit that the difference comes from the nature and diversity of images we used (narrow surrealist images compared to more various natural images), but further investigation remain needed to better understand this surprising result.
@@
Qualitatively, while the model demonstrated a robust ability to regenerate the broad category of the original images, the fidelity of content reconstruction within these images presented a more varied outcome. Some reconstructions were visually semantically close to the remembered images. However, others diverged more significantly, indicating a variability in the model’s ability to precisely capture and reconstruct the nuanced details of each imagined scene. We provide uncurated reconstructions in Fig.5 in Appendix. This variance in reconstruction fidelity can be attributed, in part, to the limitations imposed by the dataset’s size and the duration of the fMRI scanning sessions. With only 6 hours of scans and 1200 images constituting the dataset, the goal for training a model from scratch to achieve perfect reconstruction is out of reach. Nevertheless, these initial results are promising, suggesting a capacity for accurate content reproduction. It is reasonable to anticipate that with an expanded dataset, the fidelity of the reconstructed images to their original counterparts would see important improvement, enhancing both the precision and consistency of the content reproduction in future iterations.